### v1

# Self-Filtered Distillation with LLMs-generated Trust Indicators for Reliable Patent Classification

Abstract
Large language models (LLMs) increasingly generate natural language rationales to enhance interpretability, but these often contain logical errors, label mismatches, and domain-specific misalignments. Directly using such rationales as supervision risks propagating noise and undermining training stability. To address this challenge, we introduce Self-Filtered Distillation, a framework specifically tailored for patent classification, which treats LLM-generated rationales as trust signals rather than ground-truth supervision. The framework employs selective distillation guided by three unsupervised trust metrics: (1) Self-Consistency, which measures the stability of LLM-generated rationales across multiple generations; (2) Class Entailment Alignment, which assesses semantic coherence with patent-specific class definitions; and (3) LLM Agreement Scoring, which validates rationale–label plausibility. These metrics are integrated into a unified trust score that primarily weights training samples while optionally filtering out extremely low-trust cases, enabling reasoning-aware supervision. Experiments on the USPTO-2M dataset, a widely used benchmark for patent classification, show that our method outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability, establishing a reliable paradigm for leveraging reasoning-aware trust indicators in patent analytics.
1 Introduction
Patent classification is critical for organizing large-scale patent corpora into technical classification schemes, serving as the foundation for applications such as intellectual property management, technology scouting, and innovation analysis (Jiang and Goetz, 2024; Yoo et al., 2025). Unlike general-domain texts, patents exhibit a multifaceted nature: highly specialized terminology, complex claim structures, and strategically ambiguous expressions, which pose significant challenges for automated classification. Our approach is specifically tailored to the patent domain, distinguishing it from general text classification techniques. Recent advances in large language models (LLMs) offer strong potential to address these challenges by processing massive patent datasets with improved accuracy and generating natural language rationales that enhance interpretability and provide a reasoning-based foundation for trust in classification outcomes (Lee and Hsiang, 2019; Bai et al., 2024).
Figure 1: Direct distillation vs. self-filtered distillation: Weighting and filtering rationales through trust metrics yields more stable and interpretable student models.
Beyond predictive accuracy, research in NLP has increasingly emphasized explanation-driven learning as a means of enhancing interpretability and trustworthiness. This line of work leverages human-annotated rationales as auxiliary supervision signals, demonstrating that explanation can improve both generalization and interpretability (Camburu et al., 2018; DeYoung et al., 2020). While promising, such approaches incur high annotation costs and are often limited to short texts and simple classification tasks. Consequently, they fall short when applied to the classification of multifaceted and lengthy patents, where explanation-driven supervision faces domain-specific challenges.
With the advent of LLMs, research has increasingly shifted toward leveraging model-generated rationales to strengthen learning without requiring expensive human annotation (Wei et al., 2022; Hsieh et al., 2023). While this paradigm offers scalability and interpretability, LLM-generated rationales often contain logical inconsistencies, label mismatches, and semantic misalignment with domain-specific classification schemes (Turpin et al., 2023), unsuitable for multifaceted patents. For instance, an LLM may justify categorizing a digital communication patent (H04L) based on antenna structures, or misclassify an artificial intelligence patent (G06N) as general data processing (G06F). Such errors extend beyond performance degradation: in the patent domain, they can lead to misguided technology mapping, flawed market analysis, or even intellectual property disputes. Treating all rationales as direct supervisory signals risks propagating these errors and reducing training stability. Therefore, high-stakes applications of patent classification demand a framework that can quantitatively assess rationale quality and selectively incorporate only trustworthy rationales (Lampinen et al., 2022).
To address this challenge, Self-Filtered Distillation (SFD) reinterprets LLM-generated rationales as trust indicators rather than ground-truth labels, dynamically weighting their contribution and filtering out extremely low-trust cases. We propose three unsupervised trust metrics to guide the distillation process: (1) Self-Consistency, which measures the stability of multiple rationales generated for the same input; (2) Class Entailment Alignment, which evaluates semantic coherence between rationales and class-specific definitions; and (3) LLM Agreement Scoring, which employs an external verifier LLM to assess the plausibility of rationale-label pairs. These metrics are aggregated into a unified trust score that dynamically modulates the contribution of each training instance during distillation.
Empirical evaluations on the USPTO-2M dataset demonstrate that SFD surpasses traditional label-based learning and conventional distillation methods, achieving superior performance in predictive accuracy, training stability, and interpretability. This indicates that trust-based LLM rationale modulation not only enhances classification performance but also establishes a new learning paradigm for effectively leveraging LLM-generated rationales.
The primary contributions include:
  • Trust-based distillation paradigm: which treats LLM-generated rationales as trust indicators rather than ground-truth signals, reducing instability and error propagation while turning model explanations into reliable learning resources.
  • Unsupervised trust metrics: consist of Self-Consistency, Class Entailment Alignment, and LLM Agreement Scoring, which evaluate rationale quality without human annotations and provide a scalable mechanism for weighting and filtering noisy rationales.
  • Patent-domain validation: Applied to patent classification, the SFD outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability, demonstrating its value for trustworthy patent analytics.
2 Related Work
2.1 Patent Classification
Patent classification is a central task in intellectual property management, supporting prior art search, and legal decision-making. Early studies used TF–IDF features with classifiers such as SVMs or logistic regression, with limited performance due to domain-specific terminology and high-dimensional feature spaces (Fall et al., 2003).
Deep learning brings about significant improvements: CNN- and RNN-based models capture semantic and structural patterns from claims and abstracts, outperforming methods (Li et al., 2018). Still, the challenges of long documents, technical vocabulary, and concept interdependencies make patent classification harder than general text classification. Transformer-based models open a new paradigm. Domain-adapted encoders such as PatentBERT (Lee and Hsiang, 2019) and PatentSBERTa (Bekamiri et al., 2021) outperform BERTs, while generative approaches like PatentGPT show the potential of large-scale pretraining (Bai et al., 2024). Leveraging the hierarchical structure of the Cooperative Patent Classification (CPC) also helps address the complexity of multi-label classification (Zhou et al., 2021). Large-scale datasets like USPTO-2M, with millions of patents and thousands of CPC classes, further accelerate progress but also introduce issues of computational cost and severe label imbalance (Risch and Krestel, 2020).
In summary, research has advanced from statistical models to transformer-based approaches, achieving steady gains. Yet the patent nature, including a hierarchical CPC system, long documents, domain-specific language, and label imbalance, has open challenges. For high-stakes patent applications, accuracy must be complemented by transparency and interpretability, motivating more reliable, explanation-aware frameworks.
2.2 Explanation-driven Learning
Chain-of-Thought (CoT) prompting enables LLMs to solve complex reasoning tasks by generating intermediate rationales, improving predictive performance and interpretability across arithmetic reasoning, commonsense inference, and logical deduction (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2023). This has inspired the distillation research that reasoning traces from large teacher models are transferred into smaller students (Hsieh et al., 2023). For instance, KeyPoint-CoT compresses lengthy rationales into concise key points for efficiency and faithfulness (Feng et al., 2024). Yet most distillation approaches assume uniform rationale quality, which risks propagating inconsistent or erroneous explanations. This is especially problematic in patent classification, where misaligned rationales can distort label assignments.
Explanation-driven learning more broadly uses rationales as auxiliary supervision. Datasets like e-SNLI (Camburu et al., 2018) and ERASER (DeYoung et al., 2020) show that human-written explanations improve generalization and transparency, while in-context rationales aid robustness in few-shot learning (Lampinen et al., 2022; Wiegreffe et al., 2021). However, LLM-generated rationales often suffer from incompleteness, unfaithfulness, or contradictions with labels (Turpin et al., 2023). These issues are amplified in patent classification due to multifaceted patent features including lengthy claims, technical vocabulary, and complex categorization schemes.
In summary, while CoT prompting and explanation-driven learning highlight the value of LLM rationales, their unstable quality poses risks if applied directly. Our framework addresses this by treating rationales not as deterministic supervision but as trust indicators that dynamically adjust each instance’s training weight. This preserves interpretability and generalization benefits while ensuring stability and reliability in high-stakes domains such as patent classification.
2.3 Trustworthiness and Selective Learning
Ensuring the trustworthiness of LLM predictions has become a critical issue as LLMs are increasingly used in decision-sensitive contexts. A major line of work improves reliability at the inference time. Self-Consistency aggregates multiple reasoning paths for stable outputs (Wang et al., 2023), selective prediction methods allow abstention under uncertainty (Yoshikawa and Okazaki, 2023), and self-evaluation frameworks train models to assign confidence scores (Chen et al., 2023). Uncertainty-aware fine-tuning further mitigates hallucinations and stabilizes reasoning (Krishnan et al., 2024; Lin et al., 2023). While introducing effective safeguards, these methods remain largely confined to inference-level mechanisms.
Our work instead integrates trust indicators into the training process. Rather than filtering outputs post hoc, we use rationale quality to dynamically reweight training instances. Specifically, we extend Self-Consistency as a measure of rationale stability during learning, and complement it with Class Entailment Alignment and LLM Agreement Scoring to evaluate semantic coherence and external plausibility. This selective distillation strategy makes the training procedure itself reliability-aware.
Such integration is important in domains like patent classification, where low-quality rationales can propagate errors with legal and economic impact. By embedding trust indicators into distillation, our framework advances beyond inference-level reliability toward responsible training practices, ensuring model improvements are guided by explanations that are both stable and faithful.
Figure 2: Overview of the proposed Self-Filtered Distillation(SFD) framework: It consists of three stages. STEP 1: LLM-based prediction and rationale generation, where both labels and natural language rationales are produced. STEP 2: Reasoning quality evaluation through three unsupervised trust indicators: Self-Consistency (SC), Class Entailment Alignment (CEA), and LLM Agreement Score (LAS). STEP 3: Trust-aware training control, which modulates the loss contribution using soft weighting and threshold-based filtering to suppress low-quality rationales and enhance reliable signals.
3 Methodology
The Self-Filtered Distillation (SFD) framework, shown in Figure 2, quantitatively evaluates the quality of reasoning generated by LLMs and dynamically adjusts the learning contribution according to its estimated trustworthiness. The training process consists of three stages: Step 1 with LLM-based prediction and reasoning generation, Step 2 with reasoning quality evaluation, and Step 3 for trust-aware loss modulation.
3.1 LLM-Based Prediction and Reasoning Generation
In our setting, an LLM receives an input text $x$ and is prompted to generate two outputs: (1) a predicted label or a probability distribution over labels, and (2) a natural language rationale $r$ that explains the prediction. While the predicted label is used for training, the generated rationale is not treated as a ground-truth supervision signal. Instead, it serves as auxiliary information for estimating the reliability of the sample.
Formally, let $\mathcal{Y}$ denote the set of candidate labels. The model outputs a probability vector $\hat{\mathbf{y}}\in[0,1]^{|\mathcal{Y}|}$, where each component $\hat{y}_{c}$ indicates the confidence for label $c\in\mathcal{Y}$. The true label is denoted as $\mathbf{y}$. The model is trained to minimize a label-based loss using $\hat{\mathbf{y}}$. At the inference time, thresholds or task-specific decision rules are applied to derive final predictions.
Figure 3: Overview of the three trust indicators in our Reasoning Quality Evaluation: SC ensures stability, CEA verifies semantic alignment, and LAS provides external validation.
3.2 Reasoning Quality Evaluation: Lightweight Trust Scoring
Rather than serving as ground-truth supervision, LLM-generated reasoning is interpreted as a trust signal that reflects whether the reasoning is reliable enough to guide training. This enables a selective distillation mechanism: high-trust rationales contribute more to model updates, while noisy or inconsistent rationales are automatically downweighted.
To quantify the trustworthiness of rationales without relying on human annotations, we introduce three complementary unsupervised metrics, as shown in Figure 3. Each captures a distinct facet of rationale quality.
(1) Self-Consistency Scoring (SC)
To assess the logical stability of LLM-generated reasoning, we prompt the same model $k$ times with identical input $x$, obtaining a set of reasonings $\{r_{1},\dots,r_{k}\}$. The prompt used for this multi-sample generation is detailed in Appendix A. We then compute the average pairwise cosine similarity between their sentence embeddings $\phi(r_{i})$:
$\mathrm{SC}(x)=\frac{2}{k(k-1)}\sum_{i<j}\cos\left(\phi(r_{i}),\phi(r_{j})\right)$ (1)
Higher SC values indicate that the model produces consistent rationales across runs, suggesting stable reasoning with reduced stochasticity.
For instance, in the context of patent classification, when the same document concerning image classification produces divergent rationales across multiple generations, such as one emphasizing “machine learning techniques” and another focusing on “neural network architectures,” the resulting SC score is low. Such inconsistency indicates a lack of reasoning stability, which our framework interprets as reduced reliability during training.
(2) Class Entailment Alignment (CEA)
CEA measures how well the reasoning $r$ semantically aligns with each predicted label $\hat{y}$. For every label, we generate the task ontology using an LLM to obtain its definition $d_{\hat{y}}$, and then compute the cosine similarity between the embeddings of $r$ and $d_{\hat{y}}$.
[…]
By averaging over the predicted label set $\hat{Y}$, CEA evaluates the degree to which the rationale is consistent with the label space. The Cooperative Patent Classification (CPC) system follows a hierarchical structure, where broader technical domains (e.g., G06F for computer technology) contain more specific subclasses (e.g., G06N for artificial intelligence). In CEA, we explicitly leverage this property by extracting the formal definitions of predicted CPC codes through an LLM and aligning them with the generated rationale. For multi-label cases, where a patent may belong to several subclasses simultaneously, CEA averages the alignment across all candidate CPC definitions, ensuring that the rationale remains coherent with the domain-specific taxonomy.
For instance, in patent classification, when a document is assigned to H04L (digital communication) but the generated rationale emphasizes aspects of “antenna structure” rather than “signal transmission protocols,” the semantic mismatch results in a low CEA score. This reflects misalignment between the rationale and the formal CPC definition, which our framework treats as an indicator of reduced trustworthiness.
(3) LLM Agreement Scoring (LAS)
[…]
Here, $\sigma$ is the sigmoid function mapping the raw score into $[0,1]$. By introducing an external verifier, LAS provides an additional validation layer that counteracts model overconfidence and complements the other trust metrics by evaluating plausibility from an independent perspective.
For instance, in cases where the original text concerns “data encryption methods” but the generated rationale incorrectly frames the invention as addressing “data compression,” the external verifier assigns a low LAS score. This penalization captures the lack of plausibility in the rationale–label alignment and highlights the necessity of incorporating independent validation to mitigate overconfidence in model-generated explanations.
SC, CEA, and LAS offer complementary perspectives on reasoning quality: SC captures the stability across multiple generations, CEA leverages the CPC taxonomy for semantic alignment, and LAS evaluates the external plausibility through independent validation. Unlike reference-based metrics such as BLEU or ROUGE, which rely on observable textual similarity, our trust indicators assess deeper aspects of reasoning that are crucial for patent classification. This design not only supports rigorous and trustworthy selective distillation but also enables the detection of unreliable rationales that traditional similarity metrics would overlook.
3.3 Trust-Aware Training Control
The Combined Trust Score (CTS) derived from our multi-aspect evaluation is used to modulate the learning contribution of each training sample. Our framework adopts a soft weighting strategy where all samples participate in training, but their loss contributions are scaled proportionally to their estimated trustworthiness. For instance, a sample with a low trust score, $\mathrm{CTS}(x)\approx 0.5$, has its influence significantly reduced. This approach mitigates the negative impact of low-quality reasoning without entirely discarding potentially informative signals.
Formally, let the model $f_{\theta}$ predict a probability vector $\hat{\mathbf{y}}=f_{\theta}(x)\in[0,1]^{|\mathcal{Y}|}$ for an input $x\in\mathcal{X}$, with the ground-truth label denoted as $y\in\mathcal{Y}$. The task-specific loss function, $\mathrm{Loss}(\cdot)$ (e.g., cross-entropy for classification), is scaled by the trust score $\mathrm{CTS}(x)\in[0,1]$.
The final trust score, referred to as CTS, is computed as the direct average of the three previously defined metrics (SC, CEA, and LAS). This weighting scheme treats each dimension of trustworthiness as equally important, providing a balanced and robust measure of rationale quality without introducing additional learnable parameters. The CTS is thus defined as:
$\mathrm{CTS}(x)=\frac{1}{3}\big(\mathrm{SC}(x)+\mathrm{CEA}(x)+\mathrm{LAS}(x)\big).$ (4)
The overall training objective is to minimize the trust-weighted loss across all samples in data $\mathcal{D}$:
$\mathcal{L}=\sum_{x\in\mathcal{D}}\mathrm{CTS}(x)\cdot\mathrm{Loss}\big(f_{\theta}(x),\,y\big).$ (5)
In addition to this soft weighting, we introduce a thresholding mechanism to completely exclude samples with extremely low trust scores. An indicator function, ${1}\{\cdot\}$, is applied to filter out any sample whose trust score falls below a predefined threshold $\tau$:
$\mathcal{L}=\sum_{x\in\mathcal{D}}{1}\{\mathrm{CTS}(x)\geq\tau\}\cdot\mathrm{CTS}(x)\cdot\mathrm{Loss}\big(f_{\theta}(x),\,y\big).$ (6)
Candidate thresholds, such as $\tau\in\{0.1,0.2,0.3,\ldots\}$, are evaluated on a validation set, and the optimal value $\tau^{\ast}$ is selected for the final training. This strategy of combining soft weighting and hard thresholding ensures the model robust against noisy rationales while effectively leveraging high-quality reasoning signals.
4 Experiments and Results
We evaluate the proposed framework by addressing four research questions that collectively assess its effectiveness, robustness, and interpretability in patent classification.
RQ1: How does SFD compare with baseline methods? RQ2: Does threshold-based filtering improve robustness compared to soft weighting? RQ3: What is the effect of removing each trust metric? RQ4: Are trust scores correlated with human judgments of rationale quality?
4.1 Dataset
In this study, we employ the USPTO-2M dataset (United States Patent and Trademark Office, ). USPTO-2M is a large-scale dataset released by the United States Patent and Trademark Office (USPTO), consisting of more than two million patent documents and thousands of CPC codes. This dataset has become one of the most widely used benchmarks for multi-label patent classification and has been consistently adopted as a reliable evaluation resource in numerous research efforts (Lee and Hsiang, 2019; Bekamiri et al., 2021; Zhou et al., 2021).
Patent documents differ substantially from general-domain texts due to their use of specialized terminology, complex sentence structures, and strong interdependencies among technical concepts. These characteristics introduce significant challenges for automated classification. To ensure both coverage and feasibility, we restrict the label space to the subclass level of the CPC hierarchy. This setting prevents over-fragmentation of labels while maintaining sufficient representativeness of the data, as widely adopted in prior patent classification studies (Lee and Hsiang, 2019; Haghighian Roudsari et al., 2022; Risch et al., 2020).
Therefore, USPTO-2M serves as an appropriate evaluation basis for this work. It not only reflects the inherent challenges of patent data such as large scale, complexity, and hierarchical structure, but also provides a robust foundation for fair comparison across different modeling approaches.
4.2 Experimental Settings
Qwen3-30B is used as the teacher model with Qwen3-0.6B as the student model in our distillation framework. For self-consistency, we generate three rationales per instance, balancing efficiency and stability. Additional rationales yield marginal gains with higher cost. Qwen3-32B serves as the Judge LLM for LAS computation, providing an independent evaluation of rationale–label plausibility. The trust coefficients $\lambda_{1},\lambda_{2},\lambda_{3}$ are fixed at $1/3$, treating each metric equally as in the definition of $\mathrm{CTS}(x)$. We optimize the student model using the Adam optimizer with a learning rate of 1e-5. All experiments are conducted on the USPTO-2M dataset, as described in Section 4.1.
4.3 RQ1. How does SFD compare with baseline methods?
Method F1-Micro F1-Macro
Label-only Distillation 0.821 0.401
Naive CoT Distillation 0.942 0.401
Self-Consistency Distillation 0.932 0.354
LLM-as-a-Judge Distillation 0.893 0.355
Rationale-Augmented Distillation 0.944 0.369
SFD 0.960 0.404
Table 1: Performance comparison of SFD against baseline methods on USPTO-2M. Results show improvements in both Micro-F1 and Macro-F1, indicating enhanced effectiveness and robustness.
We first examine whether the proposed SFD framework outperforms existing baseline approaches. Table 1 presents the experimental results on the USPTO-2M dataset.
The label-only distillation baseline shows the weakest performance with an F1-Micro of 0.821 and an F1-Macro of 0.401. Naive CoT distillation achieves a strong F1-Micro of 0.942, but its F1-Macro remains at 0.401, indicating limited improvement on minority classes. Self-consistency distillation and LLM-as-a-judge distillation record 0.932/0.354 and 0.893/0.355 respectively, showing partial gains but lacking balance across labels. Rationale-augmented distillation reaches 0.944/0.369, which is relatively stable yet still below the best result.
By contrast, SFD achieves the top scores with 0.960 and 0.404 (F1-Micro, F1-Macro), surpassing all baselines. The improvement in Macro-F1 highlights that the method delivers reliable reasoning signals for rare categories. These findings demonstrate that treating LLM-generated rationales as trust indicators and selectively incorporating them into training yields both performance gains and improved robustness in patent classification.
4.4 RQ2. Does threshold-based filtering improve robustness compared to soft weighting?
Threshold ($\tau$) F1-Micro F1-Macro Subset Acc.
No Filtering 0.821 0.401 0.648
$\tau>0.5$ 0.820 0.400 0.646
$\tau>0.6$ 0.828 0.421 0.660
$\tau>0.7$ 0.837 0.439 0.677
$\tau>0.8$ 0.890 0.462 0.793
$\tau>0.9$ 0.960 0.404 0.930
Table 2: Results of threshold-based filtering on USPTO-2M, showing the trade-off between Micro-F1, Macro-F1, and Subset Accuracy across different $\tau$ values.
We compare two strategies: (i) soft weighting, where all samples contribute proportionally to their trust scores ($\tau=0$); and (ii) threshold-based filtering, which excludes samples with $\mathrm{CTS}(x)<\tau$. As shown in Table 2, mild filtering ($\tau>0.5$) yields limited improvement, while progressively stronger thresholds lead to more robust performance. At $\tau>0.8$, the model achieves the highest Macro-F1 (0.462), showing strong reliability for rare classes. At $\tau>0.9$, both Micro-F1 (0.960) and Subset Accuracy (0.930) reach their peak values, indicating that retaining only the most trustworthy rationales produces the most stable and reliable predictions overall. This demonstrates that high-threshold filtering effectively suppresses noisy reasoning while preserving informative signals. Consequently, we select $\tau>0.9$ as the optimal configuration for subsequent experiments.
4.5 RQ3. What is the effect of removing each trust metric?
Method F1-micro F1-macro Subset ACC.
SC only 0.932 0.354 0.898
CEA only 0.952 0.367 0.922
LAS only 0.893 0.355 0.844
w/o LAS 0.922 0.346 0.859
w/o CEA 0.797 0.262 0.695
w/o SC 0.957 0.359 0.930
$\mathrm{CTS}(x)$ 0.960 0.404 0.930
Table 3: Ablation study of trust metrics on USPTO-2M ($\tau>0.9$), reporting Micro-F1, Macro-F1, and Subset Accuracy. SC = Self-Consistency, CEA = Class Entailment Alignment, LAS = LLM Agreement Scoring.
We conduct an ablation study on the USPTO-2M dataset to evaluate the relative contributions of each trust metric. Performance is measured using F1-Micro, F1-Macro, and Subset Accuracy. Following the results of RQ2, we fix the threshold at $\tau>0.9$ to ensure consistency across experiments.
[…]
These findings suggest that CEA provides the strongest domain-grounded signal, while SC and LAS complement CEA by enhancing stability and plausibility. Their incorporation is essential not only for maximizing predictive accuracy but also for ensuring that reasoning remains interpretable and trustworthy in patent classification tasks.
4.6 RQ4. Are trust scores correlated with human judgments of reasoning quality?
Metric Pearson ($\rho$) $\Delta$ Rel. $\Delta$
SC only 0.623 - 0.062 - 9.05%
CEA only 0.612 - 0.073 - 10.66%
LAS only 0.528 - 0.157 - 22.92%
w/o LAS 0.584 - 0.101 - 14.75%
w/o CEA 0.475 - 0.210 - 30.66%
w/o SC 0.631 - 0.054 - 7.88%
$\mathrm{CTS}(x)$ 0.685 - -
Table 4: Pearson correlation ($\rho$) between trust scores and human judgments on 1,000 samples of USPTO-2M. $\Delta$ and Rel. $\Delta$ denote absolute and relative differences from $\mathrm{CTS}(x)$.
We sample 1,000 instances from USPTO-2M with varied trust scores. Three annotators rate each rationale on a 1–5 Likert scale for logical consistency, task alignment, and plausibility, with Krippendorff’s $\alpha=0.641$. Guidelines are in Appendix D.
Table 4 shows that the $\mathrm{CTS}(x)$ correlates strongest with human ratings, $\rho=0.685$. CEA is most critical, with its removal causing the largest drop at -30.66%, followed by LAS at -22.92%, and SC at -7.88%. $\mathrm{CTS}(x)$ enhances the interpretability in patent classification.
Overall, these findings demonstrate that all three metrics contribute meaningfully, with CEA playing the most critical role. $\mathrm{CTS}(x)$ provides a more reliable and interpretable signal of reasoning quality, aligning closely with human judgments in complex domains such as patent classification.
5 Conclusion
This paper proposes SFD, a framework that improves reliability and accuracy in LLM-based patent classification by leveraging reasoning quality as a supervision signal. Unlike prior methods relying solely on model outputs, our approach evaluates rationales with trust metrics and incorporates them into training. Three trust metrics comprise Stability Consistency, Class Entailment Alignment, and Logic Agreement Score. Their aggregated scores dynamically weight or filter training samples, allowing the model to emphasize high-quality rationales while suppressing noisy ones.
SFD leads to threefold contributions: (1) establishing a protocol for assessing rationale trustworthiness, (2) designing a trust-driven weighting and filtering strategy, and (3) demonstrating consistent gains across LLM backbones and patent benchmarks. Experimental results show that trust-aware supervision not only enhances predictive performance and interpretability, but also establishes a new paradigm for reliable LLM-based reasoning in high-stakes patent analytics. As for future work, we will extend the SFD framework to other high-stakes domains such as biomedical and legal texts.
Limitations
While the proposed Self-Filtered Distillation framework demonstrates consistent improvements over strong baselines, several considerations and avenues for future work remain. Our experiments are conducted on the USPTO-2M dataset, which consists of English-language patents and serves as the most widely used and reliable benchmark in patent NLP research. This choice ensures comparability and validity, yet relying solely on English patents limits the linguistic diversity of our evaluation. Future work could extend the framework to multilingual patent corpora, thereby enhancing its generality and applicability across different linguistic contexts.
In addition, the framework has been designed and validated within the patent domain. Patents exhibit typical characteristics of high-stakes text, including specialized terminology, complex structures, and strict accuracy requirements, which make them an appropriate testbed. Nevertheless, extending the framework to other high-stakes domains such as medicine, law, or scientific reporting would further demonstrate its robustness and position it as a more broadly applicable trust-aware learning paradigm.
Ethical Considerations
This study is conducted on the publicly available USPTO-2M dataset, and thus does not involve issues related to copyright or personal data. However, the proposed trust-based learning framework is designed for high-stakes domains such as patents, where misinterpretation of model-generated or model-evaluated reasoning could potentially influence legal or industrial decision-making. Therefore, the methods presented in this work are intended for research purposes, and any practical deployment should be accompanied by careful expert validation.
While the framework employs reasonable trust indicators, including Self-Consistency, Class Entailment Alignment, and LLM Agreement Scoring, these measures cannot fully replace human judgment.
For transparency, we note that ChatGPT was used to improve grammar and clarity of expression.
Appendix A Prompt for Self-Consistency (SC)
Listing 1: Prompt used to generate multiple rationales for self-consistency (SC).
prompt_template = """
You are an expert in patent classification with deep knowledge of the Cooperative Patent Classification (CPC) system. Your task is to analyze the provided patent text (abstract or representative claim) and identify the relevant CPC technical subclasses. You must also provide a detailed, logical reasoning process explaining why each subclass was selected, ensuring alignment with the CPC class definitions.
Instructions:
1. Identify one or more CPC subclasses (e.g., G06F’, H04L’, A61B’) that best represent the technical content of the patent text. Ensure the subclasses are valid and specific to at least the 4-character CPC code level.
2. Provide a rigorous reasoning process that justifies your classification decisions. The reasoning must be less than 60 words and must be coherent, grounded in the technical content of the patent, and aligned with the CPC subclass definitions.
3. Consider that a patent may belong to multiple CPC subclasses due to its multi-label nature.
4. Output your response strictly in the JSON format below, with no additional text, comments, or Markdown formatting.
Example output format:
"predicted_labels": ["B25B"],
"reasoning": "The patent text describes a mechanical hand tool, specifically a wrench. Key features include adjustable jaws for gripping nuts and bolts and a handle for applying torque, aligning with CPC subclass B25B."
Patent text:
{patent_text}
Appendix B Prompt for Class Entailment Alignment (CEA)
Listing 2: Prompt used for Class Entailment Alignment (CEA).
prompt_template = """
You are an expert on the Cooperative Patent Classification (CPC) system.
Your task is to provide a concise, official definition for the given CPC subclass code.
The definition should be a single, clear sentence. Do not add any extra explanation or introductory text.
CPC Code: {cpc_code}
Definition: """
Appendix C Prompt for LLM Agreement Scoring (LAS)
Listing 3: Prompt used for LLM Agreement Scoring (LAS).
LAS_PROMPT_TEMPLATE_CONTINUOUS = """
You are an impartial and strict judge. Your task is to evaluate if the provided Reasoning logically and accurately supports the assignment of the given Predicted Labels based on the Original Text’.
Rate the quality of the reasoning on a scale of 1 to 5.
- 1: Completely illogical, irrelevant, or hallucinatory.
- 2: Poorly reasoned, contains significant flaws.
- 3: Partially correct but has logical gaps or is not well-supported.
- 4: Mostly logical and relevant, with minor flaws.
- 5: Perfectly logical, coherent, and directly justifies the labels based on the text.
Your answer MUST be a single digit from 1 to 5.
Original Text:
{text}
Predicted Labels: {labels}
Reasoning:
{reasoning}
Based on your evaluation, what is the score for this reasoning? Answer with a single digit (1-5).
Score:"""
Appendix D Annotation Guidelines
To systematically evaluate the quality of LLM-generated rationales, we recruited three annotators and provided them with detailed guidelines to ensure consistency. Each rationale was evaluated on a 1–5 Likert scale according to three criteria: Logical Consistency, Task Alignment, and Plausibility. Below we describe the definitions, scoring criteria, and illustrative examples for each criterion.
D.1 Logical Consistency
Definition: Whether the rationale maintains internal consistency without contradictions.
Scoring scale:
  • 1 (Very Low): Contains severe contradictions that undermine the entire explanation.
    • Example: States “this invention is implemented in hardware circuits” but later claims “it operates purely as software.”
  • 2 (Low): Multiple inconsistencies, illogical flow, difficult to follow.
    • Example: Key concepts appear disconnected and the reasoning is incoherent.
  • 3 (Medium): Minor inconsistencies or ambiguities, but overall reasoning is still acceptable.
    • Example: Early part emphasizes classification algorithms, later shifts focus to preprocessing, creating mild inconsistency.
  • 4 (High): Mostly consistent with only small logical leaps or minor off-topic additions.
    • Example: Explanation is coherent around CNN-based classification but ends with a tangential remark.
  • 5 (Very High): Completely consistent, clear logical flow throughout.
    • Example: Clearly explains CNN feature extraction and softmax classification without contradictions.
D.2 Task Alignment
Definition: The extent to which the rationale aligns with the predicted CPC code definitions.
Scoring scale:
  • 1 (Very Low): Almost entirely unrelated to the CPC code.
    • Example: For H04L (digital communication), the rationale discusses “pharmaceutical compositions.”
  • 2 (Low): Largely misaligned, focusing on irrelevant aspects.
    • Example: For H04L, the rationale only describes “antenna structure” without addressing communication protocols.
  • 3 (Medium): Generally aligned but missing key elements or overly focused on secondary aspects.
    • Example: For H04L, discusses signal processing but omits transmission protocol.
  • 4 (High): Well aligned, with only minor irrelevant details.
    • Example: For H04L, primarily covers protocols and error correction, with slight mention of antenna design.
  • 5 (Very High): Perfectly aligned with CPC definitions and core technical aspects.
    • Example: For H04L, explains “data transmission protocols and error correction methods.”
D.3 Plausibility
Definition: Whether the rationale is technically factual and reasonable within the patent domain.
Scoring scale:
  • 1 (Very Low): Clearly unrealistic or factually incorrect.
    • Example: Claims “images are classified using tokenization techniques.”
  • 2 (Low): Implausible or largely inaccurate.
    • Example: “An encryption patent drastically improves signal transmission speed.”
  • 3 (Medium): Exaggerated or incomplete, but partially plausible.
    • Example: “The neural network achieves nearly perfect accuracy in classification.”
  • 4 (High): Mostly factual with minor inaccuracies.
    • Example: “CNN extracts image features and applies additional rule-based filtering.”
  • 5 (Very High): Fully factual, technically sound, and domain-appropriate.
    • Example: “CNN extracts features and softmax computes class probabilities.”
D.4 Annotation Procedure
  • Each annotator independently evaluated 500 randomly assigned samples.
  • For each rationale, scores were assigned for Logical Consistency, Task Alignment, and Plausibility, each on a 1–5 scale.
  • The final human score for each sample was computed as the average across annotators.
  • Inter-annotator reliability was measured with Krippendorff’s alpha, yielding $\alpha=0.641$, indicating acceptable agreement.