### v3

# Self-Filtered Distillation with LLMs-generated Trust Indicators for Reliable Patent Classification

Abstract
Large language models (LLMs) increasingly generate natural language rationales to enhance interpretability, but these often contain logical errors, label mismatches, and domain-specific misalignments. Directly using such rationales as supervision risks propagating noise and undermining training stability. To address this challenge, we introduce Self-Filtered Distillation, a framework tailored for patent classification that treats LLM-generated rationales as trust signals rather than ground-truth supervision. The framework employs selective distillation guided by three unsupervised trust metrics: (1) Self-Consistency, which measures the stability of LLM-generated rationales across multiple generations; (2) Class Entailment Alignment, which assesses semantic coherence with patent-specific class definitions; and (3) LLM Agreement Scoring, which validates rationale–label plausibility. These metrics are integrated into a unified trust score that primarily weights training samples while optionally filtering out extremely low-trust cases, enabling reasoning-aware supervision. Experiments on the USPTO-2M dataset show that our method consistently outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability across diverse student architectures, establishing a reliable paradigm for leveraging reasoning-aware trust indicators in patent analytics.
1 Introduction
Patent classification is critical for organizing large-scale patent corpora into technical classification schemes, serving as the foundation for applications such as intellectual property management, technology scouting, and innovation analysis (jiang2024natural; yoo2025patentmind). Unlike general-domain texts, patents exhibit a multifaceted nature: highly specialized terminology, complex claim structures, and strategically ambiguous expressions, which pose significant challenges for automated classification. Our approach is specifically tailored to the patent domain, distinguishing it from general text classification techniques. Recent advances in large language models (LLMs) offer strong potential to address these challenges by processing massive patent datasets with improved accuracy and generating natural language rationales that enhance interpretability and provide a reasoning-based foundation for trust in classification outcomes (lee2019patentbert; bai2024patentgpt).
Figure 1: Overview of the Self-Filtered Distillation (SFD) strategy. The framework evaluates LLM-generated rationales using three trust indicators. Rationales with high trust scores are incorporated into the student model training via soft weighting proportional to their trust score, while those falling below the threshold $\tau$ are filtered out to maintain training stability.
Beyond predictive accuracy, research in NLP has increasingly emphasized explanation-driven learning as a means of enhancing interpretability and trustworthiness. This line of work leverages human-annotated rationales as auxiliary supervision signals, demonstrating that explanations can improve both generalization and interpretability (camburu2018esnli; deyoung2020eraser). While promising, such approaches incur high annotation costs and are often limited to short texts and simple classification tasks. Consequently, they fall short when applied to the classification of multifaceted and lengthy patents, where explanation-driven supervision faces domain-specific challenges.
With the advent of LLMs, research has shifted toward leveraging model-generated rationales to strengthen learning without requiring human annotation (wei2022chain; hsieh2023distilling). While this paradigm offers scalability and interpretability, LLM-generated rationales often contain logical inconsistencies, label mismatches, and semantic misalignment with domain-specific classification schemes (turpin2023language) that are unsuitable for patents. For instance, an LLM may justify categorizing a digital communication patent (H04L) based on antenna structures or misclassify an artificial intelligence patent (G06N) as general data processing (G06F). Such errors extend beyond performance degradation: in the patent domain, they can lead to misguided technology mapping, flawed market analysis, or even intellectual property disputes. Treating all rationales as direct supervisory signals risks propagating these errors and reducing training stability. Therefore, high-stakes applications of patent classification demand a framework that can quantitatively assess rationale quality and selectively incorporate only trustworthy rationales (lampinen2022can).
To address this challenge, Self-Filtered Distillation (SFD) reinterprets LLM-generated rationales as trust indicators rather than ground-truth labels, dynamically weighting their contribution and filtering out extremely low-trust cases. We propose three unsupervised trust metrics to guide the distillation process: (1) Self-Consistency, which measures the stability of multiple rationales generated for the same input; (2) Class Entailment Alignment, which evaluates semantic coherence between rationales and class-specific definitions; and (3) LLM Agreement Scoring, which employs an external verifier LLM to assess the plausibility of rationale-label pairs. These metrics are aggregated into a unified trust score that dynamically modulates the contribution of each training instance during distillation.
Empirical evaluations on the USPTO-2M dataset demonstrate that SFD consistently surpasses traditional label-based learning and conventional distillation methods across diverse student architectures, achieving superior performance in predictive accuracy, training stability, and interpretability. These results indicate that incorporating unsupervised trust indicators as quality filters effectively enhances classification performance and establishes a statistically robust learning paradigm for leveraging LLM-generated rationales in complex, high-stakes domains.
The primary contributions include:
  • Trust-based distillation paradigm: treats LLM-generated rationales as trust indicators rather than ground-truth signals, reducing instability and error propagation while turning explanations into reliable learning resources.
  • Unsupervised trust metrics: consist of Self-Consistency, Class Entailment Alignment, and LLM Agreement Scoring, which evaluate rationale quality without human annotations and provide a scalable mechanism for weighting and filtering noisy rationales.
  • Patent-domain validation: Applied to patent classification, SFD outperforms label-based learning and conventional distillation in accuracy, stability, and interpretability, demonstrating value for trustworthy patent analytics.
2 Related Work
2.1 Patent Classification
Patent classification is a central task in intellectual property management, supporting prior art search, and legal decision-making. Early studies used TF–IDF features with classifiers such as SVMs or logistic regression, with limited performance due to domain-specific terminology and high-dimensional feature spaces (fall2003automated).
Deep learning brings about significant improvements: CNN- and RNN-based models capture semantic and structural patterns from claims and abstracts, outperforming methods (li2018deeppatent). Still, the challenges of long documents, technical vocabulary, and concept interdependencies make patent classification harder than general text classification. Transformer-based models open a new paradigm. Domain-adapted encoders such as PatentBERT (lee2019patentbert) and PatentSBERTa (bekamiri2021patentsberta) outperform BERTs, while generative approaches like PatentGPT show the potential of large-scale pretraining (bai2024patentgpt). Leveraging the hierarchical structure of the Cooperative Patent Classification (CPC) also helps address the complexity of multi-label classification (zhou2021hierarchical). Large-scale datasets like USPTO-2M, with millions of patents and thousands of CPC classes, further accelerate progress but also introduce issues of computational cost and severe label imbalance (risch2020patentmatch).
In summary, research has advanced from statistical models to transformer-based approaches, achieving steady gains. Yet the patent nature, including a hierarchical CPC system, long documents, domain-specific language, and label imbalance, has open challenges. For high-stakes patent applications, accuracy must be complemented by transparency and interpretability, motivating more reliable, explanation-aware frameworks.
2.2 Explanation-driven Learning
Chain-of-Thought (CoT) prompting enhances LLM performance and interpretability in complex reasoning tasks by generating intermediate rationales (wei2022chain; kojima2022zero; wang2023selfconsistency). This has inspired distillation methods where student models learn from teacher reasoning traces (hsieh2023distilling). While approaches like KeyPoint-CoT improve efficiency (feng2024keypointcot), most assume uniform rationale quality. This risks propagating errors, a critical issue in patent classification where misaligned rationales can distort label assignments. Explanation-driven learning utilizes rationales as auxiliary supervision, with human-annotated datasets demonstrating improved generalization and robustness (camburu2018esnli; deyoung2020eraser; lampinen2022can; wiegreffe2021measuring). Although metrics like ROSCOE exist to assess faithfulness (golawn2022roscoe; hase2020evaluating), LLM rationales often suffer from incompleteness and label inconsistencies (turpin2023language; ye2022unreliability; zhu2025rationales). These flaws are amplified in patent classification due to its complex technical vocabulary and categorization schemes.
While CoT and explanation-driven learning highlight the utility of rationales, their instability poses risks if applied directly. Our framework addresses this by repurposing evaluation concepts, such as consistency and alignment, as trust indicators to dynamically adjust training weights. This approach preserves interpretability benefits while ensuring stability and reliability in high-stakes domains like patent classification.
2.3 Trustworthiness and Selective Learning
Ensuring LLM trustworthiness is critical in decision-sensitive contexts. Existing work largely focuses on inference-time reliability through methods like Self-Consistency (wang2023selfconsistency) and semantic uncertainty quantification (kuhn2023semantic). Other approaches employ selective prediction or self-evaluation for confidence scoring (yoshikawa2023selective; chen2023selfeval). While uncertainty-aware fine-tuning helps mitigate hallucinations (krishnan2024enhancing; lin2023uncertainty; damani2025beyond), these safeguards remain confined to inference or require specific tuning objectives. Research on self-improvement explores bootstrapping via synthetic data, often leveraging Reinforcement Learning (RL) (zelikman2022star; huang2023selfimprove). However, these powerful methods typically rely on ground-truth verification or computationally expensive reward models to filter outputs.
Our work integrates trust indicators into the training process as a more efficient alternative. Instead of post-hoc filtering or RL loops, we dynamically reweight training instances based on rationale quality using Self-Consistency, Class Entailment Alignment, and LLM Agreement Scoring. This selective distillation strategy enables reliability-aware training without expensive annotations. In high-stakes domains like patent classification, this prevents error propagation from low-quality rationales, advancing beyond inference-level reliability toward responsible, stability-guided training.
3 Methodology
Figure 2: Overview of the proposed Self-Filtered Distillation(SFD) framework: It consists of three stages. STEP 1: LLM-based prediction and rationale generation, where both labels and natural language rationales are produced. STEP 2: Reasoning quality evaluation through three unsupervised trust indicators: Self-Consistency (SC), Class Entailment Alignment (CEA), and LLM Agreement Score (LAS). STEP 3: Trust-aware training control, which modulates the loss contribution using soft weighting and threshold-based filtering to suppress low-quality rationales and enhance reliable signals.
The Self-Filtered Distillation (SFD) framework, shown in Figure 2, quantitatively evaluates the quality of reasoning generated by LLMs and dynamically adjusts the learning contribution according to its estimated trustworthiness. The training process consists of three stages: Step 1 with LLM-based prediction and reasoning generation, Step 2 with reasoning quality evaluation, and Step 3 for trust-aware loss modulation.
3.1 LLM-Based Prediction and Reasoning Generation
In our setting, an LLM receives an input text $x$ and is prompted to generate two outputs: (1) a predicted label or a probability distribution over labels, and (2) a natural language rationale $r$ that explains the prediction. While the predicted label is used for training, the generated rationale is not treated as a ground-truth supervision signal. Instead, it serves as auxiliary information for estimating the reliability of the sample.
Formally, let $\mathcal{Y}$ denote the set of candidate labels. The model outputs a probability vector $\hat{\mathbf{y}}\in[0,1]^{|\mathcal{Y}|}$, where each component $\hat{y}_{c}$ indicates the confidence for label $c\in\mathcal{Y}$. The true label is denoted as $\mathbf{y}$. The model is trained to minimize a label-based loss using $\hat{\mathbf{y}}$. At the inference time, thresholds or task-specific decision rules are applied to derive final predictions.
Figure 3: Overview of the three trust indicators in our Reasoning Quality Evaluation: SC ensures stability, CEA verifies semantic alignment, and LAS provides external validation.
3.2 Reasoning Quality Evaluation: Lightweight Trust Scoring
Rather than serving as ground-truth supervision, LLM-generated reasoning is interpreted as a trust signal that reflects whether the reasoning is reliable enough to guide training. This enables a selective distillation mechanism: high-trust rationales contribute more to model updates, while noisy or inconsistent rationales are automatically downweighted.
To quantify the trustworthiness of rationales without relying on human annotations, we introduce three complementary unsupervised metrics, as shown in Figure 3. Each captures a distinct facet of rationale quality.
(1) Self-Consistency Scoring (SC)
To assess the logical stability of LLM-generated reasoning, we prompt the same model $k$ times with identical input $x$, obtaining a set of reasonings $\{r_{1},\dots,r_{k}\}$. The prompt used for this multi-sample generation is detailed in Appendix A. We then compute the average pairwise cosine similarity between their sentence embeddings $\phi(r_{i})$:
$\mathrm{SC}(x)=\frac{2}{k(k-1)}\sum_{i<j}\cos\left(\phi(r_{i}),\phi(r_{j})\right)$ (1)
Higher SC values indicate that the model produces consistent rationales across runs, suggesting stable reasoning with reduced stochasticity.
For instance, in the context of patent classification, when the same document concerning image classification produces divergent rationales across multiple generations, such as one emphasizing “machine learning techniques” and another focusing on “neural network architectures,” the resulting SC score is low. Such inconsistency indicates a lack of reasoning stability, which our framework interprets as reduced reliability during training.
(2) Class Entailment Alignment (CEA)
CEA measures how well the reasoning $r$ semantically aligns with each predicted label $\hat{y}$. For every label, we generate the task ontology using an LLM to obtain its definition $d_{\hat{y}}$, and then compute the cosine similarity between the embeddings of $r$ and $d_{\hat{y}}$.
[…]
By averaging over the predicted label set $\hat{Y}$, CEA evaluates the rationale’s consistency with the label space. The Cooperative Patent Classification (CPC) system follows a hierarchical structure, where broader technical domains (e.g., G06F for computer technology) contain specific subclasses (e.g., G06N for artificial intelligence). CEA leverages this by extracting definitions of predicted CPC codes via an LLM and aligning them with the rationale. For multi-label cases, CEA averages alignment across all candidate definitions, ensuring coherence with the domain taxonomy. For instance, when a document is assigned to H04L (digital communication) but the rationale emphasizes “antenna structure” rather than “signal transmission protocols,” the semantic mismatch results in a low CEA score, indicating reduced trustworthiness.
(3) LLM Agreement Scoring (LAS)
[…]
Here, $\sigma$ is the sigmoid function mapping the raw score into $[0,1]$. By introducing an external verifier, LAS provides an additional validation layer that counteracts model overconfidence and complements the other trust metrics by evaluating plausibility from an independent perspective.
For instance, in cases where the original text concerns “data encryption methods” but the generated rationale incorrectly frames the invention as addressing “data compression,” the external verifier assigns a low LAS score. This penalization captures the lack of plausibility in the rationale–label alignment and highlights the necessity of incorporating independent validation to mitigate overconfidence in model-generated explanations.
SC, CEA, and LAS offer complementary perspectives on reasoning quality: SC captures the stability across multiple generations, CEA leverages the CPC taxonomy for semantic alignment, and LAS evaluates the external plausibility through independent validation. Unlike reference-based metrics such as BLEU or ROUGE, which rely on observable textual similarity, our trust indicators assess deeper aspects of reasoning that are crucial for patent classification. This design not only supports rigorous and trustworthy selective distillation but also enables the detection of unreliable rationales that traditional similarity metrics would overlook.
3.3 Trust-Aware Training Control
The Combined Trust Score (CTS) derived from our multi-aspect evaluation is used to modulate the learning contribution of each training sample. Our framework adopts a soft weighting strategy where all samples participate in training, but their loss contributions are scaled proportionally to their estimated trustworthiness. For instance, a sample with a low trust score, $\mathrm{CTS}(x)\approx 0.5$, has its influence significantly reduced. This approach mitigates the negative impact of low-quality reasoning without entirely discarding potentially informative signals.
Formally, let the model $f_{\theta}$ predict a probability vector $\hat{\mathbf{y}}=f_{\theta}(x)\in[0,1]^{|\mathcal{Y}|}$ for an input $x\in\mathcal{X}$, with the ground-truth label denoted as $y\in\mathcal{Y}$. The task-specific loss function, $\mathrm{Loss}(\cdot)$ (e.g., cross-entropy for classification), is scaled by the trust score $\mathrm{CTS}(x)\in[0,1]$.
The final trust score, referred to as CTS, is computed as the direct average of the three previously defined metrics (SC, CEA, and LAS). This weighting scheme treats each dimension of trustworthiness as equally important, providing a balanced and robust measure of rationale quality without introducing additional learnable parameters. The CTS is thus defined as:
$\mathrm{CTS}(x)=\frac{1}{3}\big(\mathrm{SC}(x)+\mathrm{CEA}(x)+\mathrm{LAS}(x)\big).$ (4)
The overall training objective is to minimize the trust-weighted loss across all samples in data $\mathcal{D}$:
$\mathcal{L}=\sum_{x\in\mathcal{D}}\mathrm{CTS}(x)\cdot\mathrm{Loss}\big(f_{\theta}(x),\,y\big).$ (5)
In addition to this soft weighting, we introduce a thresholding mechanism to completely exclude samples with extremely low trust scores. An indicator function, ${1}\{\cdot\}$, is applied to filter out any sample whose trust score falls below a predefined threshold $\tau$:
$\mathcal{L}=\sum_{x\in\mathcal{D}}\mathbf{1}\{\mathrm{CTS}(x)\geq\tau\}\cdot\mathrm{CTS}(x)\cdot\mathrm{Loss}\big(f_{\theta}(x),y\big)$
Candidate thresholds, such as $\tau\in\{0.1,0.2,0.3,\ldots\}$, are evaluated on a validation set, and the optimal value $\tau^{\ast}$ is selected for the final training. This strategy of combining soft weighting and hard thresholding ensures the model robust against noisy rationales while effectively leveraging high-quality reasoning signals.
4 Experiments and Results
We evaluate the proposed framework by addressing five research questions that collectively assess its effectiveness, robustness, and interpretability in patent classification. RQ1: How does SFD compare with baseline methods? RQ2: Does threshold-based filtering improve robustness compared to soft weighting? RQ3: What is the effect of removing each trust metric? RQ4 : Is the proposed framework model-agnostic across different student architectures? RQ5: Are trust scores correlated with human judgments
4.1 Dataset
We employ the USPTO-2M dataset (uspto2m), a predominant benchmark in multi-label patent classification comprising over two million documents. This extensive scale encompasses diverse technical descriptions across various fields, effectively mitigating selection bias and ensuring statistically robust evaluation. Patent texts exhibit unique challenges, including specialized terminology, complex structures, and concept interdependencies, which distinguish them from general-domain texts. To balance coverage with computational feasibility, we restrict the label space to the subclass level of the Cooperative Patent Classification (CPC) hierarchy. This configuration prevents label over-fragmentation while preserving sufficient data representativeness, a standard practice in prior studies (lee2019patentbert; patentnet; Risch2020HierarchicalSeqGen). USPTO-2M serves as a robust and rigorous foundation for evaluating model performance against these inherent domain challenges.
4.2 Experimental Settings
We employ Qwen3-30B as the teacher and validate our framework across four diverse student models: Qwen3-0.6B, Phi-4-Mini, Llama-3.2-3B, and Falcon3-3B. Rationales are generated via Nucleus sampling ($T=0.6,p=0.9$), with results averaged over five independent runs to ensure reliability. For trust evaluation, we sample three rationales per instance and set equal coefficients ($\lambda=1/3$). We utilize Qwen3-32B as the judge for its strong legal reasoning capabilities and all-MiniLM-L6-v2 for embeddings. Training uses the Adam optimizer with a learning rate of 1e-5. All code and checkpoints will be released upon acceptance.
4.3 RQ1. How does SFD compare with baseline methods?
Method F1-Micro F1-Macro
Baseline Methods
Label-only Distillation 0.821 0.401
Naive CoT Distillation 0.942 0.401
Self-Consistency Distillation 0.932 0.354
LLM-as-a-Judge Distillation 0.893 0.355
Rationale-Augmented Distillation 0.944 0.369
Our Framework (SFD)
SFD w/ Qwen3-0.6B 0.960 0.404
SFD w/ Phi-4-Mini 0.969 0.503
SFD w/ Llama-3.2-3B 0.988 0.515
SFD w/ Falcon3-3B 0.988 0.556
Table 1: Performance comparison on USPTO-2M. Our SFD framework consistently outperforms strong baselines, achieving up to a 38.7% relative improvement.
Table 1 compares the performance of SFD against various distillation baselines on the USPTO-2M dataset. A critical observation is that SFD consistently outperforms the strongest baselines across all tested configurations. To rigorously validate the framework’s robustness, we conducted five independent runs for all student architectures. Every model demonstrated stable performance gains with minimal variance; for instance, SFD achieved $0.960\pm 0.006$ Micro-F1 with Qwen3-0.6B and $0.988\pm 0.003$ with Llama-3.2-3B. These results confirm that our trust-aware mechanism effectively stabilizes training regardless of the underlying backbone. Furthermore, SFD achieved a remarkable breakthrough with the Falcon3-3B backbone, recording a Micro-F1 of 0.988 and a Macro-F1 of 0.556, which represents a relative improvement of up to 38.7% over the strongest baseline. These results collectively demonstrate that SFD not only enhances predictive accuracy but also ensures high reliability in complex patent classification tasks.
4.4 RQ2. Does threshold-based filtering improve robustness compared to soft weighting?
Threshold ($\tau$) F1-Micro F1-Macro Subset Acc.
No Filtering 0.821 0.401 0.648
$\tau>0.5$ 0.820 0.400 0.646
$\tau>0.6$ 0.828 0.421 0.660
$\tau>0.7$ 0.837 0.439 0.677
$\tau>0.8$ 0.890 0.462 0.793
$\tau>0.9$ 0.960 0.404 0.930
Table 2: Effect of threshold-based filtering using Qwen3-0.6B on USPTO-2M. Progressively stricter filtering improves robustness, with $\tau>0.9$ achieving a peak Subset Accuracy of 0.930.
To evaluate the impact of filtering, we use Qwen3-0.6B as the student model. We compare soft weighting, where all samples contribute proportionally, against threshold-based filtering which excludes samples with low trust scores. As shown in Table 2, mild filtering yields limited improvement, whereas progressively stronger thresholds lead to substantial performance gains. Notably, setting the threshold at $\tau>0.9$ achieves peak performance for the Qwen3-0.6B model, recording 0.960 in Micro-F1 and 0.404 in Macro-F1, along with a significant boost in Subset Accuracy to 0.930. This indicates that retaining only the most trustworthy rationales effectively suppresses noisy reasoning while preserving informative signals. Consequently, we select $\tau>0.9$ as the optimal configuration for subsequent experiments.
4.5 RQ3. What is the effect of removing each trust metric?
Method F1-micro F1-macro Subset ACC.
SC only 0.932 0.354 0.898
CEA only 0.952 0.367 0.922
LAS only 0.893 0.355 0.844
w/o LAS 0.922 0.346 0.859
w/o CEA 0.797 0.262 0.695
w/o SC 0.957 0.359 0.930
$\mathrm{CTS}(x)$ 0.960 0.404 0.930
Table 3: Ablation study of trust metrics on USPTO-2M ($\tau>0.9$), reporting Micro-F1, Macro-F1, and Subset Accuracy. SC = Self-Consistency, CEA = Class Entailment Alignment, LAS = LLM Agreement Scoring.
We conducted an ablation study on the USPTO-2M dataset using Qwen3-0.6B to rigorously evaluate the contribution of each trust metric, with the threshold fixed at $\tau>0.9$. As shown in Table 3, the integrated trust score ($\mathrm{CTS}(x)$) yields the most robust performance, achieving 0.960 in Micro-F1 and 0.404 in Macro-F1. A hierarchical analysis of the metrics reveals distinct roles for each component. Removing CEA results in the most severe degradation, causing Subset Accuracy to plummet to 0.695. This confirms that enforcing semantic alignment with CPC definitions is the foundational requirement for correct classification. The exclusion of LAS leads to a significant drop in overall precision, indicating its critical function in filtering out plausible but hallucinated rationales. Interestingly, removing SC has a marginal impact on Micro-F1 (0.957) but notably affects Macro-F1 (0.359). This suggests that while self-consistency is less critical for clear-cut cases, it plays a vital role in stabilizing predictions for complex or minority classes. Therefore, while CEA provides the necessary domain grounding, the complementary integration of LAS and SC is essential for maximizing both reliability and generalization.
4.6 RQ4. Is the proposed framework model-agnostic across different student architectures?
Student Model Micro-F1 Macro-F1 Subset Acc.
Qwen3-0.6B 0.9600 0.4040 0.9330
Phi-4-Mini 0.9685 0.5025 0.9453
Llama-3.2-3B 0.9880 0.5153 0.9765
Falcon3-3B 0.9881 0.5556 0.9766
Table 4: Performance comparison across different student architectures using the SFD framework.
To verify whether the effectiveness of SFD is independent of specific model structures, we evaluated the framework across a diverse set of student architectures, including Qwen3, Phi-4, Llama-3.2, and Falcon3. As shown in Table 4, SFD consistently delivers robust performance across all tested backbones, confirming that the benefits of trust-aware distillation are not limited to a single model type. All models achieved high accuracy, with Micro-F1 scores exceeding 0.960 across the board. Notably, the Falcon3-3B model achieved the highest performance, with a Micro-F1 0.988 and a Macro-F1 0.556. These results demonstrate that our framework effectively transfers reasoning capabilities to various architectures, proving its generalizability and model-agnostic nature.
4.7 RQ5. Are trust scores correlated with human judgments of reasoning quality?
Metric Pearson ($\rho$) $\Delta$ Rel. $\Delta$
SC only 0.623 - 0.062 - 9.05%
CEA only 0.612 - 0.073 - 10.66%
LAS only 0.528 - 0.157 - 22.92%
w/o LAS 0.584 - 0.101 - 14.75%
w/o CEA 0.475 - 0.210 - 30.66%
w/o SC 0.631 - 0.054 - 7.88%
$\mathrm{CTS}(x)$ 0.685 - -
Table 5: Pearson correlation ($\rho$) between trust scores and human judgments on 1,000 samples of USPTO-2M. $\Delta$ and Rel. $\Delta$ denote absolute and relative differences.
We sampled 1,000 instances from USPTO-2M with varied trust scores. Three annotators rated each rationale on a 1–5 Likert scale for logical consistency, task alignment, and plausibility, with Krippendorff’s $\alpha=0.641$. Guidelines are in Appendix D. Table 5 shows that the $\mathrm{CTS}(x)$ correlates strongest with human ratings, $\rho=0.685$. CEA is most critical, with its removal causing the largest drop at -30.66%, followed by LAS at -22.92%, and SC at -7.88%. $\mathrm{CTS}(x)$ enhances interpretability in patent classification. Overall, these findings demonstrate that all three metrics contribute meaningfully, with CEA playing the most critical role. Crucially, the dominance of CEA, which anchors evaluation to objective ontology definitions, combined with the high correlation with human experts ($\rho=0.685$), confirms that our trust scoring is robust against potential model-family preference bias. Consequently, $\mathrm{CTS}(x)$ provides a reliable and interpretable signal of reasoning quality, aligning closely with human judgments in complex domains such as patent classification.
5 Conclusion
This study introduces Self-Filtered Distillation (SFD), a framework utilizing three unsupervised trust metrics to mitigate noise in LLM-generated rationales. Empirical evaluations on the USPTO-2M dataset show that SFD significantly outperforms conventional distillation, achieving up to a 38.7% relative improvement in Macro-F1. We confirmed the framework’s statistical robustness through five independent runs and its model-agnostic nature across diverse student architectures. The strong correlation with human judgments ($\rho=0.685$) validates that our trust indicators effectively capture reasoning quality. Ultimately, SFD establishes a reliable paradigm for transferring LLM reasoning to specialized domains like patent classification, ensuring both high performance and stability.
Limitations
While the proposed Self-Filtered Distillation framework demonstrates consistent improvements over strong baselines, several considerations and avenues for future work remain. Our experiments are conducted on the USPTO-2M dataset, which consists of English-language patents and serves as the most widely used and reliable benchmark in patent NLP research. This choice ensures comparability and validity, yet relying solely on English patents limits the linguistic diversity of our evaluation. Future work could extend the framework to multilingual patent corpora, thereby enhancing its generality and applicability across different linguistic contexts.
In addition, the framework has been designed and validated within the patent domain. Patents exhibit typical characteristics of high-stakes text, including specialized terminology, complex structures, and strict accuracy requirements, which make them an appropriate testbed. Nevertheless, extending the framework to other high-stakes domains such as medicine, law, or scientific reporting would further demonstrate its robustness and position it as a more broadly applicable trust-aware learning paradigm.
Ethical Considerations
This study is conducted on the publicly available USPTO-2M dataset and thus does not involve issues related to copyright or personal data. However, the proposed trust-based learning framework is designed for high-stakes domains such as patents, where misinterpretation of model-generated or model-evaluated reasoning could potentially influence legal or industrial decision-making. Therefore, the methods presented in this work are intended for research purposes, and any practical deployment should be accompanied by careful expert validation.
While the framework employs reasonable trust indicators, including Self-Consistency, Class Entailment Alignment, and LLM Agreement Scoring, these measures cannot fully replace human judgment.
For transparency, we note that ChatGPT and Gemini was used to improve grammar, clarity of expression and help to generated figures.
Appendix A Prompt for Self-Consistency (SC)
Listing 1: Prompt used to generate multiple rationales for self-consistency (SC).
prompt_template = """
You are an expert in patent classification with deep knowledge of the Cooperative Patent Classification (CPC) system. Your task is to analyze the provided patent text (abstract or representative claim) and identify the relevant CPC technical subclasses. You must also provide a detailed, logical reasoning process explaining why each subclass was selected, ensuring alignment with the CPC class definitions.
Instructions:
1. Identify one or more CPC subclasses (e.g., G06F’, H04L’, A61B’) that best represent the technical content of the patent text. Ensure the subclasses are valid and specific to at least the 4-character CPC code level.
2. Provide a rigorous reasoning process that justifies your classification decisions. The reasoning must be less than 60 words and must be coherent, grounded in the technical content of the patent, and aligned with the CPC subclass definitions.
3. Consider that a patent may belong to multiple CPC subclasses due to its multi-label nature.
4. Output your response strictly in the JSON format below, with no additional text, comments, or Markdown formatting.
Example output format:
"predicted_labels": ["B25B"],
"reasoning": "The patent text describes a mechanical hand tool, specifically a wrench. Key features include adjustable jaws for gripping nuts and bolts and a handle for applying torque, aligning with CPC subclass B25B."
Patent text:
{patent_text}
Appendix B Prompt for Class Entailment Alignment (CEA)
Listing 2: Prompt used for Class Entailment Alignment (CEA).
prompt_template = """
You are an expert on the Cooperative Patent Classification (CPC) system.
Your task is to provide a concise, official definition for the given CPC subclass code.
The definition should be a single, clear sentence. Do not add any extra explanation or introductory text.
CPC Code: {cpc_code}
Definition: """
Appendix C Prompt for LLM Agreement Scoring (LAS)
Listing 3: Prompt used for LLM Agreement Scoring (LAS).
LAS_PROMPT_TEMPLATE_CONTINUOUS = """
You are an impartial and strict judge. Your task is to evaluate if the provided Reasoning logically and accurately supports the assignment of the given Predicted Labels based on the Original Text’.
Rate the quality of the reasoning on a scale of 1 to 5.
- 1: Completely illogical, irrelevant, or hallucinatory.
- 2: Poorly reasoned, contains significant flaws.
- 3: Partially correct but has logical gaps or is not well-supported.
- 4: Mostly logical and relevant, with minor flaws.
- 5: Perfectly logical, coherent, and directly justifies the labels based on the text.
Your answer MUST be a single digit from 1 to 5.
Original Text:
{text}
Predicted Labels: {labels}
Reasoning:
{reasoning}
Based on your evaluation, what is the score for this reasoning? Answer with a single digit (1-5).
Score:"""
Appendix D Annotation Guidelines
To systematically evaluate the quality of LLM-generated rationales, we recruited three annotators and provided them with detailed guidelines to ensure consistency. Each rationale was evaluated on a 1–5 Likert scale according to three criteria: Logical Consistency, Task Alignment, and Plausibility. Below we describe the definitions, scoring criteria, and illustrative examples for each criterion.
D.1 Logical Consistency
Definition: Whether the rationale maintains internal consistency without contradictions.
Scoring scale:
  • 1 (Very Low): Contains severe contradictions that undermine the entire explanation.
    • Example: States “this invention is implemented in hardware circuits” but later claims “it operates purely as software.”
  • 2 (Low): Multiple inconsistencies, illogical flow, difficult to follow.
    • Example: Key concepts appear disconnected and the reasoning is incoherent.
  • 3 (Medium): Minor inconsistencies or ambiguities, but overall reasoning is still acceptable.
    • Example: Early part emphasizes classification algorithms, later shifts focus to preprocessing, creating mild inconsistency.
  • 4 (High): Mostly consistent with only small logical leaps or minor off-topic additions.
    • Example: Explanation is coherent around CNN-based classification but ends with a tangential remark.
  • 5 (Very High): Completely consistent, clear logical flow throughout.
    • Example: Clearly explains CNN feature extraction and softmax classification without contradictions.
D.2 Task Alignment
Definition: The extent to which the rationale aligns with the predicted CPC code definitions.
Scoring scale:
  • 1 (Very Low): Almost entirely unrelated to the CPC code.
    • Example: For H04L (digital communication), the rationale discusses “pharmaceutical compositions.”
  • 2 (Low): Largely misaligned, focusing on irrelevant aspects.
    • Example: For H04L, the rationale only describes “antenna structure” without addressing communication protocols.
  • 3 (Medium): Generally aligned but missing key elements or overly focused on secondary aspects.
    • Example: For H04L, discusses signal processing but omits transmission protocol.
  • 4 (High): Well aligned, with only minor irrelevant details.
    • Example: For H04L, primarily covers protocols and error correction, with slight mention of antenna design.
  • 5 (Very High): Perfectly aligned with CPC definitions and core technical aspects.
    • Example: For H04L, explains “data transmission protocols and error correction methods.”
D.3 Plausibility
Definition: Whether the rationale is technically factual and reasonable within the patent domain.
Scoring scale:
  • 1 (Very Low): Clearly unrealistic or factually incorrect.
    • Example: Claims “images are classified using tokenization techniques.”
  • 2 (Low): Implausible or largely inaccurate.
    • Example: “An encryption patent drastically improves signal transmission speed.”
  • 3 (Medium): Exaggerated or incomplete, but partially plausible.
    • Example: “The neural network achieves nearly perfect accuracy in classification.”
  • 4 (High): Mostly factual with minor inaccuracies.
    • Example: “CNN extracts image features and applies additional rule-based filtering.”
  • 5 (Very High): Fully factual, technically sound, and domain-appropriate.
    • Example: “CNN extracts features and softmax computes class probabilities.”
D.4 Annotation Procedure
  • Each annotator independently evaluated 500 randomly assigned samples.
  • For each rationale, scores were assigned for Logical Consistency, Task Alignment, and Plausibility, each on a 1–5 scale.
  • The final human score for each sample was computed as the average across annotators.
  • Inter-annotator reliability was measured with Krippendorff’s alpha, yielding $\alpha=0.641$, indicating acceptable agreement.